The change in entropy of a mass m of a solid substance which has a latent heat of fusion L and melts at a temperature T is

a.LT/m

b.mLln(T)

c.mLT

d.mL/T

e.L/mT
Answer #1

change of entropy s= Q/ T , here Q= mL ,

so s=mL/T , where m = is the mass of the substance , L is latent heat and T is the temperature